# The Moon as a possible source for Earth’s co-orbital bodies

### v1

@@ Abstract
There is a growing number of Earth’s co-orbital bodies being discovered. At least five of them are known to be temporarily in quasi-satellite orbits. One of those, 469219 Kamo’oalewa, was identified as possibly having the same composition as the Moon. We explore the conditions necessary for lunar ejecta to evolve into Earth’s co-orbital bodies, with particular attention to the formation of quasi-satellite orbits. We systematically investigate the parameter space of ejection velocity and geographic launch location across the entire lunar surface. The study employs numerical simulations of the four-body problem (Sun-Earth-Moon-particle) with automated classification methodology for identifying all co-orbital states. Particles are ejected from randomly distributed points covering the entire lunar surface with velocities ranging from 1.0 to 2.6 times the Moon’s escape velocity. Trajectories co-orbital to Earth are found to be a common outcome, with approximately 6.68% of all simulated particles evolving into Earth co-orbital motion and 1.92% specifically exhibiting quasi-satellite behavior. We identify an optimal ejection velocity (1.2vesc) for quasi-satellite production, yielding over 6% conversion efficiency at this specific velocity. The spatial distribution of successful ejections shows a strong preference for the equatorial regions of the trailing hemisphere. Collisions with Earth or the Moon occur for only 4% of the sample. Our extended integrations reveal exceptionally long-lived configurations, including tadpole orbits persisting for 10,000 years and horseshoe co-orbitals maintaining stability for 5,000 years. Our results strengthen the plausibility of lunar origin for Earth’s co-orbital bodies, including quasi-satellites like Kamo’oalewa and 2024PT5. We identify both ”prompt” and ”delayed” co-orbital formation mechanisms, with a steady-state production regime that could explain the presence of lunar-derived objects in Earth’s co-orbital regions despite the infrequent occurrence of major lunar impacts capable of launching meter-scale fragments.
@@
There are several distinct solutions for this co-orbital motion, including the tadpole and the horseshoe orbits (Murray and Dermott, 1999; Namouni et al., 1999). Within the class of co-orbital motions there are also the retrograde or quasi-satellite orbits. These trajectories are similar to that of satellites, but lie far out of the Hill sphere of the secondary body, and they are unstable in the inner Solar System. Compound co-orbital trajectories, including one or more of these classes of motion, with transitions between classes along the time are also possible (Wiegert et al., 1997). In general, the quasi-satellite behaviour occurs as a compound class of motion with the horseshoe motion.
@@
With the expressive number of known Near-Earth objects (NEO), almost 37 thousand bodies as of Feb. 2025111https://cneos.jpl.nasa.gov/stats/ totals.html, it is expected that NEOs constitute the main source of Earth’s co-orbital bodies. Yet, the 40 meters wide object (469219) Kamo’oalewa, currently in the most stable known quasi-satellite motion with the Earth (de la Fuente Marcos and de la Fuente Marcos, 2016), shows an L-type reflectance spectrum that matches lunar silicate material. This points towards a possible origin of the object as a fragment ejected from the Moon (Sharkey et al., 2021). More recently, the 10 meters diameter object 2024PT5 has also been identified as potentially having lunar origin based on its spectral properties (Kareta et al., 2025). This object had a close approach with Earth in September/November 2024 with a remarkably low velocity relative to Earth (de la Fuente Marcos et al., 2024).
@@
The Chinese space agency’s Tianwen-2 mission (Zhang et al., 2021) is scheduled to launch in May 2025. One of its goals is to land on the asteroid Kamo’oalewa and collect a 100 g sample that will be returned to Earth by a capsule. In addition to collecting material, remote sensing observations will take place in orbit around Kamo’oalewa. The mission results will provide critical data to test the lunar origin hypothesis and enhance our understanding of these unique objects (Venigalla et al., 2019).
@@ 2 Numerical simulations
Unlike previous studies that focused primarily on equatorial regions or limited launch sites, our test particles were ejected from randomly distributed points covering the entire lunar surface. The velocity vectors were primarily normal to the surface and outward-pointing. We also tested a control sample with non-normal velocity vectors varying randomly in azimuth while maintaining the outward direction, confirming that launch azimuth does not significantly affect outcome statistics.
Our simulations systematically explored the velocity parameter space from 1.0 up to 2.6 times the lunar escape velocity (2.4 km/s), with an ensemble of 6,000 particles for each chosen velocity, totalizing 54,000 particles. This range of velocities was selected based on preliminary results showing a significant drop in co-orbital production at higher ejection speeds, as shown in Figure 1. To assess the independence of results from the launch epoch, we conducted four separate simulation sets with different initial positions of the Moon in its orbit around Earth (at quarter-period intervals).
The system was integrated for 1,000 years with the high order integrator IAS15 available in the Rebound package (Rein and Spiegel, 2015). For selected cases, we extended the integration time to 50,000 years with the addition of Venus, Mars, and Jupiter to capture long-term stability characteristics.
We implemented automated, quantitative criteria for classifying orbital states, eliminating the subjective visual inspection methods. Specifically, we consider a particle as co-orbital when its semi-major axis remains between 0.99 and 1.01 au for at least 8 consecutive years. Quasi-satellite classification is applied when the difference between the mean longitude of the particle and the mean longitude of Earth () oscillates around zero with an amplitude less than for at least 10 years. Horseshoe orbits are identified when oscillates around with amplitude sufficient to encompass the and points, while tadpole orbits librate around either () or ).
@@ 3 Results
Our numerical simulations revealed that lunar ejecta can naturally evolve into various co-orbital configurations with Earth. From the complete set of 54,000 simulated particles, approximately 6.68% evolved into Earth co-orbital motion, with 1.92% specifically exhibiting quasi-satellite behavior. The fraction of particles that collided with either Earth or the Moon was 4.01% (3.46% with Earth and 0.55% with the Moon). Ejection velocity is a critical parameter influencing both the likelihood and stability characteristics of resulting co-orbital states.
Fig. 2 shows a representative example of the orbital evolution of a particle that becomes co-orbital with the Earth. Along this trajectory, the horseshoe motion appears in two opportunities ( yr and yr). Between the two stages in horseshoe motion, the particle experiences a quasi-satellite motion for a period of about one hundred years. The temporal evolution of the orbital elements (semi-major axis, eccentricity and inclination) clearly shows a chaotic behaviour. However, during co-orbital phases, the orbital elements exhibit significantly more regular behavior, with notably reduced variations in semi-major axis and more constrained eccentricity excursions. This regularity is particularly evident during the stages of horseshoe (orange plots) and quasi-satellite (green plots) motions.
@@
Our findings both complement and extend recent studies on the lunar origin of Earth co-orbitals. While Castro-Cisneros et al. (2023) observed that approximately 6% of particles launched from the lunar equator became co-orbital with Earth within 5,000 years, our comprehensive survey covering the entire lunar surface yields a comparable 6.68% co-orbital production rate, validating their results while providing a more complete spatial distribution. However, our finding that 1.92% of all particles exhibit quasi-satellite behavior is significantly higher than previously reported values. Jiao et al. (2024) found that within 10 million years after ejection from Giordano Bruno crater, the percentage of particles in Earth co-orbital states reached at most 1%, with only rare cases becoming long-lived quasi-satellites. This discrepancy likely stems from our finding of an optimal ejection velocity (1.2) for quasi-satellite production, which yields over 6% conversion efficiency at this specific velocity.
@@
The full statistical analysis of our 54,000 simulated particles yielded 6.68% of trajectories evolving into Earth co-orbital motion, with 1.92% specifically exhibiting quasi-satellite behavior. Collisions occurred in only 4.01% of cases (3.46% with Earth and 0.55% with the Moon). We found that ejection velocity strongly influences co-orbital outcomes, with particles ejected at 1.2 produced the highest proportion of quasi-satellites, exceeding 6% at this specific velocity. This optimal ejection window represents a significant refinement compared to earlier studies that reported lower co-orbital conversion efficiencies (Jiao et al., 2024).
@@
Our integration timeframes revealed exceptional orbital stability in certain cases, far exceeding previous estimates. The 1,000-year simulations identified quasi-satellites and horseshoe co-orbitals maintaining their respective states throughout the entire integration period. More significantly, our extended 50,000-year integrations discovered extremely long-lived configurations: tadpole orbits persisting for 10,000 years and horseshoe co-orbitals maintaining stability for 5,000 years. These longevity values substantially exceed the stability timeframes reported by Castro-Cisneros et al. (2023), who primarily observed persistence on timescales of hundreds of years. The significantly longer orbital stability we observed might be attributed to our approach of sampling the complete lunar surface rather than equatorial regions alone and systematically exploring ejection velocities, which revealed the critical threshold for generating stable co-orbital configurations.
@@ 4 Lunar origin of co-orbitals
Fig. 6 shows the geographical distribution of initial conditions on the lunar surface that lead to Earth co-orbital objects. The map uses a lunar coordinate system where longitude points toward Earth, longitude represents the leading side (the direction of lunar orbital motion), and longitude corresponds to the trailing side. The density distribution reveals a latitudinal dependence, with a higher concentration of ejections that resulted in co-orbital motion occurring near the equatorial regions (latitude ). The map was generated using particle data at all simulated velocities and was aligned to match the four different Earth-Moon-Sun geometries. This equatorial preference is observed in the trailing side of the Moon (longitude ), where the normalized density reaches its maximum values. A secondary concentration appears in the leading hemisphere (longitude ), although with lower density values. This indicates that equatorial ejections, especially those from the trailing side, tend to be more efficient at producing Earth co-orbital objects.
@@ 5 Final remarks
Our numerical simulations demonstrate that lunar ejecta can naturally evolve into Earth co-orbital configurations, though with limited frequency. From our comprehensive sample of 54,000 simulated particles, approximately 6.68% evolved into Earth co-orbital motion, with only 1.92% specifically exhibiting quasi-satellite behavior. This finding extends previous work by Castro-Cisneros et al. (2023), who reported similar overall co-orbital production rates (6%) but focused primarily on equatorial launch sites. Our study systematically explores the entire lunar surface and reveals a non-monotonic relationship between ejection velocity and co-orbital formation, with particles ejected at 1.2 showing a marked increase in quasi-satellite production efficiency (reaching 6% at this specific velocity). The spatial distribution of successful ejection sites shows a strong preference for the equatorial regions of the trailing hemisphere, consistent with theoretical expectations and previous studies. This preferential region appears particularly important for generating long-lived co-orbital configurations, as evidenced by our geographic density analysis. Our extended integrations revealed orbital stability timeframes may significantly exceed previous estimates, with tadpole orbits persisting for up to 10,000 years and horseshoe co-orbitals maintaining stability for up to 5,000 years—an order of magnitude longer than the stability periods reported by Castro-Cisneros et al. (2023), who primarily observed persistence on timescales of hundreds of years.
Our results strengthen the plausibility of lunar origin for Earth’s co-orbital bodies, including quasi-satellites like Kamo’oalewa and 2024 PT5. While the production efficiency of such bodies remains low, our identification of an optimal ejection velocity window (1.2) and the extended orbital stability of certain configurations provide a viable pathway for lunar material to populate Earth’s co-orbital zones. Unlike Jiao et al. (2024), who focused on ejecta specifically from the Giordano Bruno crater and reported co-orbital percentages of at most 1%, our broader survey reveals higher production rates (6.68% overall and up to 6% for quasi-satellites at optimal velocity) and identifies critical parameters governing co-orbital formation. Our approach also differs from Castro-Cisneros et al. (2025), who examined the influence of Earth’s eccentricity on co-orbital outcomes while maintaining equatorial launch sites, as we systematically explore the entire parameter space of ejection velocity and geographic location.
We also identified a distinct of “prompt” and “delayed” co-orbital formation mechanisms through our temporal analysis. After an initial peak in the first 100 years, we observe a transition to a relatively constant rate of co-orbital entries (approximately 500 events per 100-year interval for short-lived states and about 75 events per interval for long-lived configurations). This steady-state production regime has not been previously reported and suggests a fundamental mechanism by which lunar material can continuously replenish Earth’s co-orbital population even centuries after the initial impact event. This mechanism could explain the presence of lunar-derived objects like Kamo’oalewa and 2024 PT5 in Earth’s co-orbital regions despite their inherently temporary orbital configurations and the infrequent occurrence of major lunar impacts. Our automated classification methodology also represents an advancement over the subjective visual inspection methods. By establishing quantitative criteria for identifying all co-orbital states (tadpole, horseshoe, and quasi-satellite), we provide a consistent, reproducible framework for population statistics that captures the full spectrum of dynamical behaviors, including transitions between states.
Future work will extend these simulations to include additional planetary perturbations and longer integration times, potentially revealing even more long-lived co-orbital configurations. Further investigation of the relationship between ejection conditions and orbital stability characteristics may provide additional constraints on the origin of specific Earth co-orbital objects and enhance our understanding of material exchange within the Earth-Moon system.

### v2

@@ Abstract
There is a growing number of Earth’s co-orbital bodies being discovered. At least five of them are known to be temporarily in quasi-satellite orbits. One of those, 469219 Kamo’oalewa, was identified as possibly having the same composition as the Moon. We explore the conditions necessary for lunar ejecta to evolve into Earth’s co-orbital bodies, with particular attention to the formation of quasi-satellite orbits. We systematically investigate the parameter space of ejection velocity and geographic launch location across the entire lunar surface. The study employs numerical simulations of the four-body problem (Sun-Earth-Moon-particle) with automated classification methodology for identifying all co-orbital states. Particles are ejected from randomly distributed points covering the entire lunar surface with velocities ranging from 1.0 to 2.6 times the Moon’s escape velocity. Trajectories co-orbital to Earth are found to be a common outcome, with approximately 6.15% of all simulated particles evolving into Earth co-orbital motion and 1.92% specifically exhibiting quasi-satellite behavior. We identify an optimal ejection velocity (1.2vesc) for quasi-satellite production, yielding over 6% conversion efficiency at this specific velocity. The spatial distribution of successful ejections shows a strong preference for the equatorial regions of the trailing hemisphere. Collisions with Earth or the Moon occur for only 4% of the sample. Our results strengthen the plausibility of lunar origin for Earth’s co-orbital bodies, including quasi-satellites like Kamo’oalewa and 2024 PT5. We identify both ”prompt” and ”delayed” co-orbital formation mechanisms, with a steady-state production regime that could explain the presence of lunar-derived objects in Earth’s co-orbital regions despite the infrequent occurrence of major lunar impacts capable of launching meter-scale fragments.
@@
There are several distinct solutions for this co-orbital motion, including the tadpole and the horseshoe orbits (Murray and Dermott, 1999; Namouni et al., 1999). Within the class of co-orbital motions there are also the retrograde or quasi-satellite orbits. These trajectories are similar to that of satellites, but lie far out of the Hill sphere of the secondary body, and they are unstable in the inner Solar System. Compound co-orbital trajectories, including one or more of these classes of motion, with transitions between classes along the time are also possible (Wiegert et al., 1997). In general, the quasi-satellite behaviour occurs as a compound class of motion with the horseshoe motion. Theoretical studies by Christou and Georgakarakos (2021) have demonstrated that deep Earth co-orbitals can survive for extended periods under specific dynamical conditions.
@@
With the significant number of known Near-Earth objects (NEO), almost 37 thousand bodies as of Feb. 2025111https://cneos.jpl.nasa.gov/stats/ totals.html, it is expected that NEOs constitute the main source of Earth’s co-orbital bodies. Yet, the 40 meters wide object (469219) Kamo’oalewa, currently in the most stable known quasi-satellite motion with the Earth (de la Fuente Marcos and de la Fuente Marcos, 2016), shows an L-type reflectance spectrum that matches lunar silicate material. This points towards a possible origin of the object as a fragment ejected from the Moon (Sharkey et al., 2021). More recently, the 10 meters diameter object 2024 PT5 was discovered (Fuentes-Muñoz et al., 2024) and has also been identified as potentially having lunar origin based on its spectral properties (Kareta et al., 2025). This object had a close approach with Earth in September/November 2024 with a remarkably low velocity relative to Earth (de la Fuente Marcos et al., 2024). Additional evidence for lunar origin comes from the minimoon 2020 CD3, which exhibited spectral properties more consistent with lunar material than typical near-Earth asteroids during its temporary capture phase (Bolin et al., 2020).
@@
The Chinese space agency’s Tianwen-2 mission (Zhang et al., 2021) launched on a Long March 3B rocket on May 28, 2025. One of its goals is to land on the asteroid Kamo’oalewa and collect a 100 g sample that will be returned to Earth by a capsule. The probe is expected to reach the asteroid in July 2026. In addition to collecting material, remote sensing observations will take place in orbit around Kamo’oalewa. The mission results will provide critical data to test the lunar origin hypothesis and enhance our understanding of these unique objects (Venigalla et al., 2019).
@@ 2 Numerical simulations
Unlike previous studies that focused primarily on equatorial regions or limited launch sites, our test particles were ejected from randomly distributed points covering the entire lunar surface using the Marsaglia (1972) algorithm for uniform sampling on a sphere. This uniform sampling ensures that our geographical distribution analysis accurately reflects the relative efficiency of different lunar regions for producing co-orbital objects, without bias toward any particular location. The velocity vectors were primarily normal to the surface and outward-pointing. We also tested a control sample with non-normal velocity vectors varying randomly in azimuth while maintaining the outward direction, confirming that launch azimuth does not significantly affect outcome statistics.
Our simulations systematically explored the velocity parameter space from 1.0 up to 2.6 times the lunar escape velocity (2.4 km/s), with an ensemble of 6,000 particles for each chosen velocity, totaling 54,000 particles. This range of velocities was selected based on preliminary results showing a significant drop in co-orbital production at higher ejection speeds, as shown in Figure 1. To assess the independence of results from the launch epoch, we conducted four separate simulation sets with different initial positions of the Moon in its orbit around Earth (at quarter-period intervals).
The system was integrated for 1,000 years with the high order integrator IAS15 available in the Rebound package (Rein and Spiegel, 2015). To verify integration quality, we monitored timestep evolution throughout sample integrations and confirmed that the adaptive timesteps selected by IAS15 maintained required numerical precision even during dynamically sensitive phases such as quasi-satellite states.
We implemented automated, quantitative criteria for classifying orbital states, eliminating the subjective visual inspection methods. Specifically, we consider a particle as co-orbital when its semi-major axis remains between 0.99 and 1.01 au for at least 10 consecutive years. This time threshold excludes transient passages through co-orbital regions that do not represent sustained dynamical states.
The quasi-satellite classification is applied when the difference between the mean longitude of the particle and the mean longitude of Earth () oscillates around zero with an amplitude less than also for at least 10 years. This amplitude limit ensures complete capture of quasi-satellite states without missing or truncating capture intervals. Horseshoe orbits are identified when oscillates around with amplitude sufficient to encompass the and points, while tadpole orbits librate around either () or ). Throughout each integration, we continuously monitored the mean longitude difference between particles and Earth at every timestep, recording entry and exit times for both quasi-satellite and general co-orbital states according to our classification criteria. This approach captured all co-orbital episodes and dynamical state transitions during the 1,000-year simulations.
@@ 3 Results
Our numerical simulations revealed that lunar ejecta can naturally evolve into various co-orbital configurations with Earth. From the complete set of 54,000 simulated particles, approximately 6.15% evolved into Earth co-orbital motion, with 1.92% specifically exhibiting quasi-satellite behavior. The fraction of particles that collided with either Earth or the Moon was 4.01% (3.46% with Earth and 0.55% with the Moon). Ejection velocity is a critical parameter influencing both the likelihood and stability characteristics of resulting co-orbital states.
Fig. 2 shows a representative example of the orbital evolution of a particle that becomes co-orbital with the Earth. Along this trajectory, the horseshoe motion appears in two time periods ( yr and yr). Between the two stages in horseshoe motion, the particle experiences a quasi-satellite motion for a period of about one hundred years. The temporal evolution of the orbital elements (semi-major axis, eccentricity and inclination) clearly shows a chaotic behaviour. However, during co-orbital phases, the orbital elements exhibit significantly more regular behavior, with notably reduced variations in semi-major axis and more constrained eccentricity excursions. This regularity is particularly evident during the stages of horseshoe (orange plots) and quasi-satellite (green plots) motions.
@@
Our findings both complement and extend recent studies on the lunar origin of Earth co-orbitals. While Castro-Cisneros et al. (2023) observed that approximately 6% of particles launched from the lunar equator became co-orbital with Earth within 5,000 years, our comprehensive survey covering the entire lunar surface yields a comparable 6.15% co-orbital production rate, validating their results while providing a more complete spatial distribution. However, our finding that 1.92% of all particles exhibit quasi-satellite behavior is significantly higher than previously reported values. Jiao et al. (2024) found that within 10 million years after ejection from Giordano Bruno crater, the percentage of particles in Earth co-orbital states reached at most 1%, with only rare cases becoming long-lived quasi-satellites. This discrepancy likely stems from our finding of an optimal ejection velocity (1.2) for quasi-satellite production, which yields over 6% conversion efficiency at this specific velocity.
@@
Some particles experience multiple co-orbital episodes during the integration, including cases where the same object exhibits both short-lived ( years) and long-lived ( years) co-orbital periods. For the classification in Fig. 4, we assign each particle to the category corresponding to its longest co-orbital duration to avoid double counting. However, to preserve information about multiple co-orbital entries and the full range of co-orbital behaviors, Fig. 5 records every individual entry event into co-orbital states, allowing us to capture the complete temporal evolution of co-orbital formation throughout the simulation.
@@
The full statistical analysis of our 54,000 simulated particles yielded 6.15% of trajectories evolving into Earth co-orbital motion, with 1.92% specifically exhibiting quasi-satellite behavior. Collisions occurred in only 4.01% of cases (3.46% with Earth and 0.55% with the Moon). We found that ejection velocity strongly influences co-orbital outcomes, with particles ejected at 1.2 produced the highest proportion of quasi-satellites, exceeding 6% at this specific velocity. This optimal ejection window represents a significant refinement compared to earlier studies that reported lower co-orbital conversion efficiencies (Jiao et al., 2024).
@@
Our integration timeframes revealed exceptional orbital stability in certain cases, far exceeding previous estimates. The 1,000-year simulations identified quasi-satellites and horseshoe co-orbitals maintaining their respective states throughout the entire integration period.
@@ 4 Lunar origin of co-orbitals
Fig. 6 shows the geographical distribution of initial conditions on the lunar surface that lead to Earth co-orbital objects. The map uses a lunar coordinate system where longitude points toward Earth, longitude represents the leading side (the direction of lunar orbital motion), and longitude corresponds to the trailing side. The geographical distribution analysis properly accounts for the spherical lunar surface geometry, with density calculations normalized per unit area to ensure accurate representation despite the rectangular projection used for visualization.
The density distribution reveals a latitudinal dependence, with a higher concentration of ejections that resulted in co-orbital motion occurring near the equatorial regions (latitude ). The map was generated using particle data at all simulated velocities and was aligned to match the four different Earth-Moon-Sun geometries. This equatorial preference is observed in the trailing side of the Moon (longitude ), where the normalized density reaches its maximum values. A secondary concentration appears in the leading hemisphere (longitude ), although with lower density values. This indicates that equatorial ejections, especially those from the trailing side, tend to be more efficient at producing Earth co-orbital objects.
@@
Our results suggest that detailed laboratory analysis of lunar meteorites or returned samples could potentially constrain their launch locations on the lunar surface. The strong geographical preferences we observe, particularly the enhanced production from equatorial regions of the trailing hemisphere, indicate that compositional and chronological data from samples could be compared against known crater distributions and ages in these favorable regions to narrow down potential source locations.
@@ 5 Final remarks
The numerical simulations demonstrate that lunar ejecta can evolve into Earth co-orbital configurations with limited frequency. Ejection velocity strongly influences co-orbital formation, with particles launched at 1.2 yielding the highest quasi-satellite production efficiency. Geographic analysis reveals preferential co-orbital formation from equatorial regions of the trailing hemisphere, consistent with theoretical expectations for lunar material capture into Earth’s co-orbital zones. These findings support the lunar origin hypothesis for Earth’s co-orbital bodies, including quasi-satellites like Kamo’oalewa and 2024 PT5. The comprehensive parameter space exploration reveals both “prompt” and “delayed” co-orbital formation mechanisms. The delayed mechanism exhibits steady-state production rates over centuries, providing a pathway for continuous replenishment of Earth’s co-orbital population between major lunar impact events. The automated classification methodology developed here provides quantitative criteria for identifying all co-orbital states (tadpole, horseshoe, and quasi-satellite), establishing a reproducible framework for population statistics that captures transitions between dynamical states.
Our results indicate that lunar ejecta may remain in Earth’s vicinity or impact our planet, with implications for lunar resource extraction activities. Excavation of lunar regolith requires mechanical forces (Wilkinson and DeGennaro, 2007) that could accelerate particles beyond the 2.38 km/s escape velocity (Just et al., 2020), particularly in the low-gravity lunar environment. International space law frameworks currently lack specific provisions for debris liability from extraterrestrial mining operations (Hubbard et al., 2024). The identified preferential ejection regions suggest that mining activities in the lunar trailing hemisphere equatorial zones present elevated risks for generating Earth-directed debris. Implementation of debris mitigation protocols will be essential for safe lunar resource extraction.
Future investigations incorporating additional planetary perturbations and extended integration times may reveal longer-lived co-orbital configurations. The relationship between ejection parameters and orbital stability merits further study to constrain the origins of specific Earth co-orbital objects. Comparative analysis of the delayed capture mechanism with resonant phenomena in other solar system contexts may yield insights into general capture dynamics in multi-body systems.
Data availability
The particle outcome data from the numerical simulations presented in this work are publicly available at https://doi.org/10.5281/zenodo.16857659.